OpenStudy (anonymous):

Solve: (r -1)(2r +3)(r + 6) < 0

OpenStudy (anonymous):

find the solutions of each factor(whats in the parenthesis)

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

there's nothing given if that was your question

OpenStudy (anonymous):

do I go ahead and distribute?

OpenStudy (anonymous):

how there is (r-1)(2r+3)(r+6), for each of those you go ahead and find the solutions...for example r-1=0 r=1 and you do that for all of them...once you have those you plug it back in to see your answer

OpenStudy (anonymous):

oh i see okay one sec

OpenStudy (anonymous):

so 1) r = 1 2) r = -3/2 3) r = -6 right?

OpenStudy (anonymous):

correct! so now plug each one of those back into the equation to see which one ends up <0

OpenStudy (anonymous):

oka i can take it from there. thank you so much!

OpenStudy (anonymous):

youre welcome :D